Suppose a ball is drawn at random from a box containing three white and three black
balls. After a ball is drawn, it is then replaced and another drawn. What is the
probability that of the first four balls drawn, exactly two are white?
1
Expert's answer
2020-06-02T19:12:29-0400

Probability of first drawn white "= \frac{3C2\times3C2}{6C2} = \frac{3}{5}"


Probability of second drawn white with replacement "= \frac{3C2\times3C2}{6C2} = \frac{3}{5}"


probability of exactly two are white = Probability of first drawn white "\times" Probability of second drawn white with replacement



probability of exactly two are white "={\frac{3}{5}}\times\frac {3}{5}= \frac {9}{25}"